# What does really matter in image goal navigation?

### v1

@@ 4.1 Benchmarking architectures
Table 1 presents a comparison of different network architectures trained with Sliding=True. We observe that the ChannelCat architecture obtains excellent performance, in SR almost approaching the levels of DEBiT-b, which, however, benefits from a more complex architecture based on cross-attention, and in particular pre-training on cross-view completion and RPVE. Without pre-training, the DEBiT architecture is not exploitable, as also reported in [8]. Both ResNet9 and ViT Late Fusion architectures underperform, which we explain by the simplicity of comparing observation and goal image late, at the outputs of the encoder $\phi$. Adding SpaceToDepth does not seem to have a significant impact neither on ChannelCat nor on Late Fusion ResNet9: being able to compare local patches across images in a single layer through convolutions does not translate into numerical gains.
[…]
@@
The high impact of the Sliding simulator setting came as a big surprise to us. While its role in sim2real transfer of point goal navigation was established in [33], the common conjecture was that it impacted “raw” navigation performance, i.e., finding free space, avoiding obstacles, short, capabilities which are independent of the extraction of goal directional information. Our experiments, however, suggest that the particular choice of binocular visual encoder $\phi$ is impacted by this hyper-parameter, which in theory should be related to motion only.
@@
(a) Load all “false” $f$$\ast$ $f$$\ast$ $f$$\ast$ $f$$\ast$ 31.7 18.7
(b) Load all “true” $t$$\ast$ $t$$\ast$ $t$$\ast$ $t$$\ast$ 54.6 27.5
(c) Load all “true” $t$$\rightarrow$ $t$$\rightarrow$ $t$$\rightarrow$ $t$$\rightarrow$ 65.7 34.1
@@
[…]
This then raises the question, whether (i) the two agents trained in their respective settings have similar capabilities but perform differently purely due to the difference in task difficulty, or (ii) whether, for some reason, training with sliding actually leads to different and perhaps better performing agent capabilities. We tested these hypotheses by performing experiments loading the weights of an agent trained with Sliding=True and finetuning it with Sliding=False. For these experiments we used the ChannelCat + SpaceToDepth agent, for which the difference in SR dropped from 83.6% to 31.7% when sliding was switched off (Tables 1 and 2).
Table 3 shows the results of these experiments. Taking the agent trained with Sliding=True and validating it on False, row (b), makes SR jump to 54.6%, compared to the baseline of 31.7%, trained on False only. This is a surprising result. Although the performance is a far cry of the 83.6% obtained by training and validating with True, it indicates that a certain amount of transfer is possible — some capabilities are learned with sliding enabled. Finetuning this agent for 100M steps on False then provides further gains and yields SR=65.7%, row (c).
To pinpoint the effect, we loaded different parts of the agent trained with Sliding=True: only the perception part (weights of the visual encoder $\phi$), only the action part (GRU $h$, policy $\pi$, encoder $\zeta$ of the previous action) or both. Transferring the action part of the agent, ie. GRU $h$, policy $\pi$, encoder $\zeta$ of the previous action, rows (d) and (e), does not lead to any discernible performance: learning the upstream visual encoder $\phi$ from scratch through pre-loaded downstream weights of the GRU and policy seems a very hard task. However, transferring the perception weights $\phi$ and training the action part from scratch, rows (f) and (g), leads to exploitable results, but significantly below loading the full agent weights. With a SR of 38.5% for the finetuning version, row (g), these are still higher than the baseline of 31.7% trained from scratch on False. In Table 4 we perform similar experiments with DEBiT which, however, uses a pre-trained visual encoder $\phi$ that is kept frozen. We load the rest of the agent trained with Sliding=True and evaluate it with False (b): in this case, no transfer happens and performance drops significantly. Finetuning the agent trained with True on False, (c), improves the Sliding=True policy, but not to the extent of reaching the performance of the agent trained on-distribution, i.e. Sliding=False, arguably because RPVE pre-training of perception is already providing a strong visual encoder.
@@
(a) Load all “false” $p$$\ast$ $f$$\ast$ $f$$\ast$ $f$$\ast$ 81.7 52.0
(b) Load all “true” $p$$\ast$ $t$$\ast$ $t$$\ast$ $t$$\ast$ 59.5 35.4
(c) Load all “true” $p$$\ast$ $t$$\rightarrow$ $t$$\rightarrow$ $t$$\rightarrow$ 79.6 46.9
@@
We attempt to draw the following lessons from these experiments. The difference between Sliding=True or Sliding=False cannot only be described by the task difficulty alone. The easier task (True) allows to learn additional capabilities, which transfer to the harder task, and which are partially related to perception (since the performance in Table 3(g) $>$ 3(a)), and also related to action / sequential decision taking, since 3(b) $\gg$ 3(a) and 3(b) $\gg$ 3(g). However, we conjecture that the transfer of the knowledge stored in the action component is also highly linked to improving the perception skills, as it does not happen when $\phi$ is pre-trained and frozen (Table 4(a) $\gg$ 4(b), and 4(a) $>$ 4(c)). One possible explanation could be that pure training with False leads to undertraining of, both, action and perception: the policy keeps getting stuck (which we empirically confirmed) and does not learn to cope with the last meters of each episode; this, in turn, leads to undertraining the comparison between the (hardly ever seen) goals and observations.
Model s2d Backbone S %corr.poses %corr.vis.
(Table nr. + row) 1m,10° 2m,20° $\mathbf{{<}0.05}$
@@
Late Fusion 2a ResNet9  7.8 26.8 13.2
Late Fusion 2b ResNet9  8.7 28.5 16.1
ChannelCat 2d ResNet9  9.8 26.9 13.8
ChannelCat 2e ResNet9 12.5 31.9 19.2
ChannelCat 3c ResNet9 $\rightarrow$ 18.2 41.4 21.1
ChannelCat 3d ResNet9 $\rightarrow$  5.8 22.9  6.7
ChannelCat 3e ResNet9 $\rightarrow$  7.2 26.1 11.9
ChannelCat 3g ResNet9 $\rightarrow$ 18.6 41.6 21.0
Cross-att 1h,2h DEBiT-b N/A 92.1 96.8 88.8
@@
Table 5: Probing relative pose and visibility estimation: representations $\phi(\mathbf{o},\mathbf{g})$ trained with an RL (navigation) loss are frozen, and then we train a probing head. The last block of methods in cursive is not comparable, as the encoders were pre-trained: DEBiT [8] was pre-trained on RPVE losses, DINOv2 [50] with SSL. We report RPVE performance on a hold-out set. s2d=SpaceToDepth, S = Sliding=True. The third block shows agents which have been finetuned ($\rightarrow$) from True to False, cf. Table 3.
4.3 Probing RPVE capabilities
Given that a certain limited capacity for visual reasoning has been learned by the visual encoder $\phi$ when the full agent has been trained for navigation with Sliding=True, we pursue this question further and directly investigate how well the different visual encoders can deal with extracting directional information. Or, motivated differently, given the potential impact on relative pose estimation of training these visual encoders from navigation losses only without pose supervision, we evaluate how well they work for this task. We took the frozen visual encoders $\phi$ of the agents evaluated in Tables 1 and 2 and trained probing heads $p$ on top of them predicting relative pose and visibility (RPVE):
In the lines of [8], relative pose between the observed image and the goal image is composed of two components, translation $\mathbf{t}{\in}\mathbb{R}^{3}$ and a rotation matrix $\mathbf{R}{\in}\mathbb{R}^{3\times 3}$. Visibility quantifies the amount of overlap between the two images, necessary since in navigation settings the goal image might not even be observed at certain moments. It is defined as the proportion of $16{\times}16$ patches of the goal image $\mathbf{g}$ which are visible in the observed image $\mathbf{o}$.
The probing heads $p$ are adapted to the different architectures, while at the same time providing a comparable capacity of $\sim{3M}$ parameters for each head variant. They collect features at the last spatial representation (feature map for ResNet, patch embeddings for ViT, concatenation of these for Late Fusion representations) and first token-wise linearly project to a lower dimension, before flattening to obtain a global representation which is passed to an MLP with 1024 hidden units. The low dimension of the projection changes according to architectures (as the spatial resolution, i.e., the number of tokens change) to make the parameter counts similar: 64 for ViTs, 192 (resp. 16) for ResNet9 with (resp. without) SpaceToDepth.
@@ 4.3 Probing RPVE capabilities
Figure 3: Nav vs. Rel-pose: navigation perf. (SR,%) plotted against pose est. probing accuracy (for err ${<}2m,20^{\circ}$) for 4 types of visual encoders $\phi$: trained w. sliding, trained w/o sliding, pre-trained w. RPVE, trained w. sliding, finetuned w/o; (LF = Late Fusion, CC=ChannelCat). The dashed line relates the finetuned models to the same model trained w/o sliding.
Figure 4: Analysis of navigation behavior: Sankey plots show the distribution of success/failure codes over 944 test episodes for different models, and their “flow” between certain pairs of models. For instance, the strength of the connection between “Time out” (left) and “success” (right) indicates how many episodes toggled from one to the other when switching from the left to the right model.
We generated a probing dataset by combining the 3D scene datasets Gibson [74], MP3D [14], and HM3D [52], following their standard train/val scene splits, as done in  [8]. We sample pairs of points randomly in the scene, calculate the shortest path between them, then extract 10 intermediate poses on the path. These 10 poses are then each combined with the end point to form pairs of observation $\mathbf{o}$ and goal $\mathbf{g}$. GT for relative pose is directly available from the simulator, and visibility can be estimated from the two point clouds generated from the respective images. We generated around 68M image pairs total.
where $i$ indexes image pairs over the probing dataset, $\mathbf{t}_{i}^{*},\mathbf{R}_{i}^{*},v_{i}^{*}$ denote ground truth values, $\mathbf{1_{.}}$ is the binary indicator function, |.||.|| . | denotes the $L_{1}$ loss and $\tau$ is a threshold.
@@
Results – Table 5 gives results comparing RPVE performance on a validation set for all tested visual encoders, and Figure 3 provides a scatter plot relating it to navigation performance (taken from Tables 1, 2, 3). While the ChannelCat architecture obtains high navigation performance in the Sliding=True setting, 3(a), its pose estimation performance remains limited. Late Fusion models generally provide low performance both for navigation and pose estimation. In the Sliding=False setting, 3(a), the navigation performance significantly drops for the ChannelCat models, as already discussed in Section 4.1, and also somewhat for pose estimation. Transferring the models trained with True on the False setting provides gains not only in navigation performance, as already discussed, but also on pose estimation — only when at least the weights of the visual encoder $\phi$ are transferred, but in particular when the whole agent is transferred and finetuned. This further corroborates our conjecture, that the perception model is undertrained when sliding is disabled. Finally, the best performance is obtained by the DEBiT model based on ViTs and cross-attention, both in navigation and pose probing. However, the pose estimation performance is not comparable for obvious reasons — the model has been pre-trained and frozen for this task. Its performance is given for information.
In Table 5 we also provide pose probing results on two additional models, where we explore whether the pose estimation performance of late fusion can be increased by using two different pre-trained and frozen encoders, namely DINOv2 and the ViT encoder part of DEBiT, without the cross-attention layers combining them [^6] DEBiT first encodes each image with a ViT and then passes both embeddings to a decoder using cross-attention, cf. [8].. They both underperform, further corroborating the importance of comparing visual representations early, allowing to compute correspondences between local parts of images.
@@ 4.4 Analysis of navigation behavior
In Figure 4 we provide Sankey plots showing the differences in navigation behavior as a distribution of success/error types over 994 test episodes evaluated in environments with Sliding=False, and how this behavior varies across different models and different training settings. Here we distinguish between unsuccessful episodes due to Time out, when the agent does not complete an episode in the 1000 steps budget, and Early stop, when the agent terminates the episode but is further away than the 1m success threshold.
Left, the ChannelCat agent trained on Sliding=False, Table 3(a), stops too early in most of the episodes, confirming that its components are not properly trained, probably due to the difficulty of the task. Next, finetuning on False the ChannelCat agent trained with True, Table 3(c), allows to obtain an agent with good navigation capabilities. It improves the ChannelCat agent trained with True, Table 3(b), in the middle in Figure 4, whose main failure mode is time out. We observe that most of these time outs occur far from the goal and are due to the agent getting stuck, likely because it learned to slide along obstacles, which is not possible with False. On the 4th column, DEBiT trained with True, Table 4(b), exhibits a similar behavior, while DEBiT trained with False, Table 4(a), on the last column, can effectively solve the complex navigation task, likely thanks to its pretrained, high capacity, perception module.

### v3

@@ 4.1 Benchmarking architectures
Tab. 1 summarizes results for different architectures and settings. With Sliding=True, ChannelCat (d)-(f) obtains excellent performance, close to DEBiT-B (h) which has a larger, more complex architecture, and is pre-trained on RPVE. Without pre-training, DEBiT is not exploitable, as also reported in [8]. Late Fusion architectures (a)-(c) underperform, and SpaceToDepth has no significant impact on either ChannelCat or Late Fusion models : being able to compare local patches across images in a single layer through convolutions does not translate into numerical gains. With Sliding=False the trends change dramatically. While the impact on the (previously already underperforming) Late Fusion architecture is similar, ChannelCat, (d)-(f), now breaks down and performance is halved, or less. In contrast, DEBiT is able to cope well with the more realistic Sliding=False setting, arguably because of its strong pre-trained visual encoder, confirming the importance of visual pre-training.
To summarize, coming back to Q1 and Q2 raised in Sec. 1, our conclusions are the following:
R1:
In realistic settings (Sliding=True), agents trained with RL alone perform significantly worse than DEBiT with pre-trained visual encoder.
R2:
Early Fusion agents perform better than Late Fusion counterparts, with early patch-wise fusion being the best option, but requiring pre-training.
Additionally, we found that recent success of agents based on small CNNs trained end-to-end with RL only, was mostly due to setting Sliding=True in Habitat.
@@
The huge impact of the Sliding simulator setting came somewhat as a surprise. While its role in sim2real transfer of point goal navigation was established in [33], the common conjecture was that it impacted “low-level navigation”, i.e., finding free space, avoiding obstacles, capabilities which are independent of the extraction of goal directional information. Our experiments, however, suggest that the choice of binocular visual encoder $\phi$ is impacted by this hyper-parameter, which should be related to motion only.
This then raises the question, whether (i) the two agents trained in their respective settings have similar capabilities but perform differently purely due to the difference in task difficulty, or (ii) whether, for some reason, training with sliding actually leads to different and perhaps better performing agent capabilities. We tested these hypotheses by performing experiments loading the weights of an agent trained with Sliding=True and finetuning it with Sliding=False. For these experiments we used the ChannelCat + SpaceToDepth agent, for which the difference in SR dropped from 83.6% to 31.7% when sliding was switched off (cf. Tab. 1).
Tab. 2 shows the results of these experiments. Taking the agent trained with Sliding=True and validating it on False, row (b), makes SR jump to 54.6%, compared to the baseline of 31.7%, trained on False only. This is a highly surprising result. Although the performance is a far cry of the 83.6% obtained by training and validating with True, it indicates that a certain amount of transfer is possible — some capabilities are learned with sliding enabled. Finetuning this agent for 100M steps on False provides further gains and yields SR=65.7%, row (c).
@@
(a) Load all “false” $\ast ff\ast$ $\ast ff\ast$ 31.7 18.7
(b) Load all “true” $\ast tt\ast$ $\ast tt\ast$ 54.6 27.5
(c) Load all “true” $\rightarrow tt\rightarrow$ $\rightarrow tt\rightarrow$ 65.7 34.1
@@
(f) Load perception ”true” $t\ast$ $circlearrowright \circlearrowright$ 26.4 14.3
(g) Load perception “true” $t\rightarrow$ $circlearrowright \circlearrowright$ 38.5 20.3
@@
To pinpoint the effect, we loaded different parts of the agent trained with Sliding=True: only the perception part (weights of the visual encoder $\phi$), only the action part (GRU $h$, policy $\pi$, previous action encoder $\zeta$) or both. Transferring the action part of the agent, i.e. GRU $h$, policy $\pi$, encoder $\zeta$, rows (d) and (e), does not lead to any discernible performance: learning the upstream visual encoder $\phi$ from scratch through pre-loaded downstream weights of the GRU and policy seems a very hard task. However, transferring the perception weights $\phi$ and training the action part from scratch, rows (f) and (g), leads to exploitable results, although significantly lower than loading the full agent weights. With a SR of 38.5% for the finetuning version, row (g), these are still higher than the baseline of 31.7% trained from scratch on False. In Tab. 3 we perform similar experiments with DEBiT which, however, uses a pre-trained visual encoder $\phi$ that is kept frozen. We load the rest of the agent trained with Sliding=True and evaluate it with False (b): in this case, no transfer happens and performance drops significantly. Finetuning the agent trained with True on False, (c), improves the Sliding=True policy, but not to the extent of reaching the performance of the agent trained on-distribution, i.e. Sliding=False, arguably because RPVE pre-training of perception is already providing a strong visual encoder.
@@
(a) Load all “false” $p\ast$ $\ast ff\ast$ $f\ast$ 81.7 52.0
(b) Load all “true” $p\ast$ $\ast tt\ast$ $t\ast$ 59.5 35.4
(c) Load all “true” $p\ast$ $\rightarrow tt\rightarrow$ $t\rightarrow$ 79.6 46.9
@@
We draw the following lessons from these experiments. The difference between Sliding=True or False cannot only be described by the task difficulty alone. The easier task (True) allows to learn additional capabilities, which transfer to the harder task, and which are partially related to perception (since the performance in Tab. 2(g) >> 2(a)), and also related to action / sequential decision taking, since Tab. 2(b) $\gg$ 2(a) and Tab. 2(b) $\gg$ 2(g). However, we conjecture that the transfer of the knowledge stored in the action component is also highly linked to improving the perception skills, as it does not happen when $\phi$ is pre-trained and frozen (Tab. 3(a) $\gg$ 3(b), and Tab. 3(a) >> 3(c)). One possible explanation is that only training with False leads to undertraining of, both, action and perception: the policy keeps getting stuck (which we empirically confirmed) and does not learn to cope with the last meters of each episode; this, in turn, leads to undertraining the comparison between the (hardly ever seen) goals and observations.
4.3 Probing RPVE capabilities
Given that a certain limited capacity for visual reasoning has been learned by the visual encoder $\phi$ when the full agent has been trained for navigation with Sliding=True, we pursue this question further and directly investigate how well the different visual encoders can deal with extracting directional information. Or, motivated differently, given the potential impact on relative pose estimation of training these visual encoders from navigation losses only without pose supervision, we evaluate how well they work for this task. We took the frozen visual encoders $\phi$ of the agents evaluated in Tab. 1 and trained probing heads $p$ on top of them predicting relative pose and visibility (RPVE):
In line with [8], relative pose between the observed image and the goal image is composed of two components, translation $\mathbf{t}{\in}\mathbb{R}^{3}$ and a rotation matrix $\mathbf{R}{\in}\mathbb{R}^{3\times 3}$. Visibility quantifies the amount of overlap between the two images, necessary since in navigation settings the goal image might not even be observed at certain moments. It is defined as the proportion of $16{\times}16$ patches of the goal image $\mathbf{g}$ which are visible in the observed image $\mathbf{o}$.
Model %corr.poses %corr.vis.
(Table nr. + row) s2d Backbone S 1m,10° 2m,20° $\mathbf{{<}0.05}$
@@ 4.3 Probing RPVE capabilities
Late Fusion 1a ResNet9  7.8 26.8 13.2
Late Fusion 1b ResNet9  8.7 28.5 16.1
ChannelCat 1d ResNet9  9.8 26.9 13.8
ChannelCat 1e ResNet9 12.5 31.9 19.2
ChannelCat 2c ResNet9 $\rightarrow$ 18.2 41.4 21.1
ChannelCat 2d ResNet9 $\rightarrow$  5.8 22.9  6.7
ChannelCat 2e ResNet9 $\rightarrow$  7.2 26.1 11.9
ChannelCat 2g ResNet9 $\rightarrow$ 18.6 41.6 21.0
Cross-att 1h DEBiT-B N/A 92.1 96.8 88.8
@@
Table 4: Probing relative pose and visibility estimation: representations $\phi(\mathbf{o},\mathbf{g})$ trained with an RL (navigation) loss are frozen, and then we train a probing head. The last block of methods in italic is not comparable, as the encoders were pre-trained: DEBiT [8] was pre-trained on RPVE losses, DINOv2 [52] with SSL. We report RPVE performance on a hold-out set. s2d=SpaceToDepth, S = Sliding=True. The third block shows agents which have been finetuned ($\rightarrow$) from True to False, cf. Tab. 2.
The probing heads $p$ are adapted to the different architectures, while at the same time providing a comparable capacity of $\sim{3M}$ parameters for each head variant. They collect features at the last spatial representation (feature map for ResNet, patch embeddings for ViT, concatenation of these for Late Fusion representations) and first linearly project token-wise to a lower dimension, before flattening to obtain a global representation which is passed to an MLP with 1024 hidden units. The low dimension of the projection changes according to architectures (as the spatial resolution, i.e., the number of tokens change) to make the parameter counts similar: 64 for ViTs, 192 (resp. 16) for ResNet9 with (resp. without) SpaceToDepth.
As in [8], we generated a probing dataset by combining the 3D scene datasets Gibson [77], MP3D [14], and HM3D [55], following their standard train/val scene splits. We sample pairs of points randomly in the scene, calculate the shortest path between them, and extract 10 intermediate poses on the path. These 10 poses are then each combined with the end point to form pairs of observation $\mathbf{o}$ and goal $\mathbf{g}$. Ground-truth for relative pose is directly available from the simulator, and visibility can be estimated from the two point clouds generated from the respective images. We generated around 68M image pairs total.
where $i$ indexes image pairs over the probing dataset, $\mathbf{t}_{i}^{*},\mathbf{R}_{i}^{*},v_{i}^{*}$ denote ground truth values, $\mathbf{1_{.}}$ is the binary indicator function, $|.|$ denotes the $L_{1}$ loss and $\tau$ is a threshold.
@@
Figure 3: Nav vs. Rel-pose: navigation perf. (SR,%) plotted against pose est. probing accuracy (% for err $2m,20^{\circ}$) for 4 types of visual encoders $\phi$: trained w. sliding, trained w/o sliding, pre-trained w. RPVE, trained w. sliding, finetuned w/o; (LF = Late Fusion, CC=ChannelCat). The dashed line relates the finetuned models to the same model trained w/o sliding.
Figure 4: Analysis of navigation behavior: Sankey plots show the distribution of success/failure codes over 994 test episodes for different models, and their “flow” between certain pairs of models. For instance, the strength of the connection between “Time out” (left) and “success” (right) indicates how many episodes toggled from one to the other when switching from the left to the right model.
@@
ResultsTab. 4 gives results comparing RPVE performance on a validation set for all tested visual encoders, and Fig. 3 provides a scatter plot relating it to navigation performance (taken from Tab. 1, 2). While the ChannelCat architecture obtains high navigation performance in the Sliding=True setting, 3(a), its pose estimation performance remains limited. Late Fusion models generally provide low performance both for navigation and pose estimation. In the Sliding=False setting, 3(a), the navigation performance significantly drops for the ChannelCat models, as already discussed in Sec. 4.1, and also somewhat for pose estimation. Transferring the models trained with True on the False setting provides gains not only in navigation performance, as already discussed, but also on pose estimation — only when at least the weights of the visual encoder $\phi$ are transferred, but in particular when the whole agent is transferred and finetuned. This further corroborates our conjecture, that the perception model is undertrained when sliding is disabled. Finally, the best performance is obtained by the DEBiT model based on ViTs and cross-attention, both in navigation and pose probing. However, the pose estimation performance is not comparable for obvious reasons — the model has been pre-trained and frozen for this task. Its performance is given for information.
In Tab. 4 we also provide pose probing results for two additional models, where we explore whether the pose estimation performance of Late Fusion can be increased by using two different pre-trained and frozen encoders, namely DINOv2 and the ViT encoder part of DEBiT, without the cross-attention layers combining them [^5] DEBiT first encodes each image with a ViT and then passes both embeddings to a decoder using cross-attention, cf. [8].. They both underperform, further corroborating the importance of comparing visual representations early, allowing to compute correspondences between local image parts. These experiments support a positive answer to Q3 in Sec. 1: there is a clear correlation between navigation and RPE performance.
@@ 4.4 Analysis of navigation behavior
In Fig. 4 we provide Sankey plots showing the differences in navigation behavior as a distribution of success/error types over 994 test episodes evaluated in environments with Sliding=False, and how this behavior varies across different models and different training settings. Here we distinguish between unsuccessful episodes due to Time out, when the agent does not complete an episode in the 1000 steps budget, and Early stop, when the agent terminates the episode but is further away than the 1m success threshold.
Left, the ChannelCat agent trained on Sliding=False, Tab. 2(a), stops too early in most of the episodes, confirming that its components are not properly trained, probably due to the difficulty of the task. Next, finetuning on False the ChannelCat agent trained with True, Tab. 2(c), allows to obtain an agent with good navigation capabilities. It improves the ChannelCat agent trained with True, Tab. 2(b), in the middle in Fig. 4, whose main failure mode is time out. We observe that most of these time-outs occur far from the goal and are due to the agent getting stuck, likely because it learned to slide along obstacles, which is not possible with False. On the 4th column, DEBiT trained with True, Tab. 3(b), exhibits a similar behavior, while DEBiT trained with False, Tab. 3(a), on the right-most column, can effectively solve the complex navigation task, arguably thanks to its pretrained, high capacity, perception module.